# EVA-Gaussian: 3D Gaussian-based Real-time Human Novel View Synthesis under Diverse Camera Settings

### v1

@@ 4.3 Gaussian Attribute Estimation
To complete the estimation of 3D Gaussian maps $\{\bm{M}_{i}\}_{i=1}^{n}$, we employ a shallow U-Net $\mathcal{D}^{A}_{\bm{\theta}_{2}}$ to estimate the remaining attributes $\bm{O}_{i}$, $\bm{S}_{i}$, $\bm{Q}_{i}$, $\bm{F}_{i}$. This network takes the estimated 3D Gaussian position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ from the first stage in Sec. 4.2 and the original RGB images $\{\bm{I}_{i}\}_{i=1}^{n}$ as input, and outputs the 3D Gaussian attributes $\bm{O}_{i}$, $\bm{S}_{i}$, $\bm{Q}_{i}$, $\bm{F}_{i}$, which is expressed as:
@@
The resulting estimated 3D Gaussian maps $\{\bm{M}_{i}\}_{i=1}^{n}=\{\bm{P}_{i},\bm{O}_{i},\bm{S}_{i},\bm{Q}_{i},\bm{F}_{i}\}_{i=1}^{n}$ are then utilized to render novel views using the process described in Sec. 3. The network $\mathcal{D}^{A}_{\bm{\theta}_{2}}$ is trained by using a combination of MSE loss and structural similarity index measure (SSIM) (Wang et al., 2004) loss between the rendered novel view image $\hat{\bm{I}}^{0}$ and the ground truth $\bm{I}^{\text{gt}}$ as follows:
@@ 4.4 Feature Splatting and Refinement
The 3D Gaussian position maps $\bm{P}_{i}$ estimated in Sec. 4.2 inevitably contain some degree of error, which may lead to distortions and artifacts in the rendered RGB images. To mitigate these issues, we propose a post-splatting refinement method to correct the position estimates. Recent studies (Berriel Martins & Civera, 2024) have demonstrated that feature vector representations can capture scene information more effectively than spherical harmonics, resulting in significant improvements in novel view synthesis, particularly in scenarios with limited overlapping views. Inspired by this finding, we incorporate a feature vector, i.e., $\bm{f}^{u}_{i}\in\mathbb{R}^{32}$ mentioned in Sec. 4.1, as an additional attribute for each Gaussian to more precisely capture its spatial characteristics.
During the splatting process, we first aggregate the 3D Gaussians from all source views. Then, the color values of these 3D Gaussians are rendered using Eq. 3. Concurrently, the feature values of the 3D Gaussians are splatted onto the image plane using a modified $\alpha$-blending function as follows:
@@
Moreover, we employ a carefully designed recurrent U-Net $\mathcal{D}_{\theta_{3}}^{R}$ that takes both the RGB and feature images as input and project them onto the RGB space for the final output through $L$ recurrent loops. This recurrent procedure is expressed as follows:
$\hat{\bm{I}}^{l}=\mathcal{D}_{\bm{\theta}_{3}}^{R}(\hat{\bm{I}}^{l-1}\oplus\bm{F}_{\text{novel}}),\hat{\bm{I}}^{l}\in\mathbb{R}^{H\times W\times 3},\bm{F}_{\text{novel}}\in\mathbb{R}^{H\times W\times 32},l\in\left\{1\cdots L\right\}.$ (11)
Similar to the Gaussian attribute estimation, the loss function for supervising the final output is a combination of MSE loss and SSIM loss between refined image $\hat{\bm{I}}^{L}$ and ground truth $\bm{I}^{\text{gt}}$ as follows:
@@
Figure 4: Attribute regularization. We regularize the opacities and scales of Gaussians, as well as the position mismatch among the Gaussians in the landmark collection. The optimization of position mismatch terminates when it falls below a specific tolerance.
@@ 4.5 Attribute Regularization
Since human faces are critical for identification and expression understanding, improving the reconstruction of human faces is much more important than that of other body parts. Previous works like GPS-Gaussian (Zheng et al., 2024) treat the entire human body equally and ignore expressive information contained in human faces. Moreover, they fail to ensure the consistency between depth maps and 3D Gaussian locations, resulting in sub-optimal reconstruction quality in the facial regions.
To address this issue, we propose a regularization term to enhance the overall reconstruction quality. Specifically, our proposed anchor loss regularizes the scales and opacities of Gaussians to ensure consistency between the geometry of predicted depth maps and the 3D Gaussian positions. It also aligns the Gaussians from different views to force their locations to the same landmark. We adopt MediaPipe (Lugaresi et al., 2019) to annotate human facial landmarks and compute the anchor loss to regularize the 3D landmark Gaussian scales, opacities, and positions as follows:
$\mathcal{L}_{\text{anchor}}=$ $\sum_{i,j\in\mathbb{V}}\sum_{m_{i}\in\mathbb{M}_{i},m_{j}\in\mathbb{M}_{j}}\max\left\{||\Pi^{-1}(\bm{m}_{i},\bm{P}_{i}(\bm{m}_{i}))-\Pi^{-1}(\bm{m}_{j},\bm{P}_{j}(\bm{m}_{j}))||_{2},t\right\}$
$+\lambda_{\text{opacity}}\sum_{i=1}^{N}||\bm{O}_{i}\text{log}(\bm{O}_{i})||_{1}+\lambda_{\text{scale}}\sum_{i=1}^{N}||\bm{S}_{i}||_{2},$ (13)
where $\{\mathbb{M}_{i}\}_{i=1}^{n}$ denotes the collection of all landmarks on the 2D image plane, $\mathbb{V}$ denotes the collection of source views, and $\Pi^{-1}$ represents the process of reprojection from 2D image to 3D space.
Since the MediaPipe landmark estimation is not perfectly accurate, we introduce a factor $t$ to control the tolerance for mismatch errors. This tolerance facilitates the optimization by activating the loss only when the landmark distance exceeds $t$. Therefore, this approach optimizes the facial reconstruction loss to a sufficiently low level and avoids being misguided by potential errors in the MediaPipe estimation. This procedure is illustrated in Fig. 4.
By integrating the loss functions in the three stages, i.e., $\mathcal{L}_{\text{depth}},\mathcal{L}_{\text{render}},\mathcal{L}_{\text{refine}}$, and the proposed regularization term $\mathcal{L}_{\text{anchor}}$, the overall training loss of the proposed EVA-Gaussian is given by:
$\mathcal{L}_{\text{EVA-Gaussian}}=\mathcal{L}_{\text{depth}}+\lambda_{1}\mathcal{L}_{\text{render}}+\lambda_{2}\mathcal{L}_{\text{refine}}+\lambda_{3}\mathcal{L}_{\text{anchor}},$ (14)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weights used to balance the different loss terms.
Since the 3D Gaussian position and attribute estimation stages can be executed within tens of milliseconds, and feature refinement is lightweight, taking less than ten milliseconds, EVA-Gaussian is capable of rapidly reconstructing 3D human subjects from a collection of RGB images and rendering novel views in a real-time manner.
@@ 5 Experiments
Table 1: Comparison with feed-forward 3D reconstruction methods at a resolution of 256$\times$256. Better results are marked in a deeper color.
$\Delta$=45° THuman2.0 THumansit inference time
@@
5.1 Experiment Setup
Implementation details. Our EVA-Gaussian is trained on 1024$\times$1024 pixel images across multiple training views using a single NVIDIA A800 GPU for 100K iterations with the AdamW (Loshchilov & Hutter, 2017) optimizer, unless otherwise specified. For the 3D Gaussian position estimation stage, it is first pretrained under the supervision of ground truth depth maps. Baselines are trained using their publicly available code.
Datasets. We conduct experiments on two open-source human body datasets: THuman2.0 (Yu et al., 2021b) and THumanSit (Zhang et al., 2023). THuman2.0 contains 526 unique human models with their corresponding SMPL parameters, among which 100 individuals are randomly selected for our evaluation. The THumanSit dataset has a similar structure, containing 72 human models with around 60 poses for each, and we randomly choose 5 individuals with all poses for our evaluation.
Metrics. We report results on commonly used metrics: PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018), computed over the entire image, as some methods may produce artifacts outside the human bounding box (Lin et al., 2022; Zheng et al., 2024). We also include the inference time to demonstrate the real-time reconstruction capability of our method.
Table 2: Comparison of feed-forward human reconstruction methods under different camera angle settings, at a resolution of $1024\times 1024$. Better results are marked in a deeper color. Notably, GPS-Gaussian fails to work effectively when $\Delta=90^{\circ}$, as it is unable to meet its rectification requirement.
@@ 5.1 Experiment Setup
Figure 5: Qualitative comparison on THuman2.0 and THumansit. EVA-Gaussian achieves superior novel view rendering quality under diverse camera settings. Additional visualization results are provided in Appendix B.
Table 3: Comparison with GPS-Gaussian under different camera number settings. The results in bold represent the best performance. Our EVA-Gaussian achieves SOTA performance across various metrics, primarily due to the multi-view consistency enabled by our proposed EVA module.
@@
w/o EVA module w/o feature refinement w/o anchor loss Full model
@@
5.2 Stereo Reconstruction
Comparison with state-of-the-art feed-forward reconstruction methods. We first compare our approach against state-of-the-art (SOTA) feed-forward reconstruction methods, including ENeRF (Lin et al., 2022), pixelSplat (Charatan et al., 2024), MVSplat (Chen et al., 2024), MVSGaussian (Liu et al., 2024b), and GPS-Gaussian (Zheng et al., 2024). All experiments are conducted in a stereo-view setting, where the angle between the two camera views $\Delta=45^{\circ}$. The attention modules in the scene reconstruction methods (Charatan et al., 2024; Liu et al., 2024b; Chen et al., 2024) are inefficient in their utilization of GPU memory, limiting their ability to train effectively at a high resolution of 1024$\times$1024. Therefore, we also conduct a fair comparison of all methods at a resolution of 256$\times$256. The quantitative results presented in Table 1 demonstrate that EVA-Gaussian achieves the best novel view quality in terms of PSNR, SSIM, and LPIPS, while maintaining the second-fastest inference speed.
@@ 5.2 Stereo Reconstruction
5.3 Multi-view Reconstruction
We conduct experiments to compare our method against GPS-Gaussian under multi-view settings. Table 3 presents the quantitative results. Our method demonstrates a significant advantage over the baseline, with a more than 1.5 dB improvement. Notably, the rendering quality of GPS-Gaussian drops significantly due to the mismatch among multiple inferences. In contrast, our method maintains high performance, thanks to the cross-view consistency ensured by our proposed EVA module.
5.4 Ablation Study
We conduct a detailed ablation study on THuman2.0 in a stereo-view setting, where the angle between the two views $\Delta=45^{\circ}$, as shown in Table 4 and Fig. 6. We gradually incorporate the EVA module, feature refinement module, and anchor loss to evaluate their individual contributions. The absence of the EVA module results in significant degradation across all metrics, and the network struggles to perform multi-view 3D Gaussian geometry prediction. When feature refinement is excluded, visualizations reveal artifacts in critical areas, such as the hands and feet. Moreover, the lack of anchor loss leads to unreliable geometry predictions, particularly in the facial region, which in turn degrades the performance across all metrics, with a notable impact on LPIPS.
@@ 6 Conclusion
In this paper, we introduce EVA-Gaussian, a novel real-time 3D human reconstruction pipeline that employs multi-view attention-based 3D Gaussian position estimation and comprehensive feature refinement. To ensure robust performance, the method is trained using both photometric loss and anchor loss. Quantitative and qualitative evaluations on benchmark datasets demonstrate that our EVA-Gaussian achieves state-of-the-art performance while maintaining a competitive inference speed, particularly under sparse camera settings.

### v2

@@ 3.3 Gaussian Attribute Estimation
To complete the estimation of 3D Gaussian maps $\{\bm{M}_{i}\}_{i=1}^{n}$, we employ a shallow U-Net $\mathcal{D}^{A}_{\bm{\theta}_{2}}$ to estimate the remaining attributes $\bm{O}_{i}$, $\bm{S}_{i}$, $\bm{Q}_{i}$, $\bm{F}_{i}$. This network takes the estimated position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ from the first stage in Sec. 3.2 and the original RGB images $\{\bm{I}_{i}\}_{i=1}^{n}$ as input, and outputs 3D Gaussian attributes $\bm{O}_{i}$, $\bm{S}_{i}$, $\bm{Q}_{i}$, $\bm{F}_{i}$, which is expressed as:
@@
The resulting estimated 3D Gaussian maps $\{\bm{M}_{i}\}_{i=1}^{n}=\{\bm{P}_{i},\bm{O}_{i},\bm{S}_{i},\bm{Q}_{i},\bm{F}_{i}\}_{i=1}^{n}$ are then utilized to render novel views through the $\alpha$-blending mechanism, as in the vanilla 3DGS [6]. The network $\mathcal{D}^{A}_{\bm{\theta}_{2}}$ is trained by using a combination of MSE loss and structural similarity index measure (SSIM) [28] loss between the rendered novel view image $\hat{\bm{I}}^{0}$ and the ground truth $\bm{I}^{\text{gt}}$ as follows:
@@ 3.4 Feature Splatting and Refinement
The 3D Gaussian position maps $\bm{P}_{i}$ estimated in Sec. 3.2 inevitably contain some degree of error, which may lead to distortions and artifacts in the rendered RGB images. To mitigate these issues, we propose a post-splatting refinement method to correct the position estimates. Recent studies [1] have demonstrated that feature vector representations can capture scene information more effectively than spherical harmonics, resulting in significant improvements in novel view synthesis, particularly in scenarios with limited overlapping views. Inspired by this finding, we incorporate a feature vector, i.e., $\bm{f}^{u}_{i}\in\mathbb{R}^{32}$ mentioned in Sec. 3.1, as an additional attribute for each Gaussian to more precisely capture its spatial characteristics.
During the splatting process, we first aggregate the 3D Gaussians from all source views. Then, the color values of these 3D Gaussians are rendered using $\alpha$-blending mechanism from 3D Gaussian Splatting [6]. Concurrently, the feature values of the 3D Gaussians are splatted onto the image plane using a modified $\alpha$-blending function:
@@
Moreover, we employ a carefully designed recurrent U-Net $\mathcal{D}_{\theta_{3}}^{R}$ that takes both the RGB and feature images as input and project them onto the RGB space for the final output through $L$ recurrent loops. This procedure is expressed as:
$\hat{\bm{I}}^{l}=\mathcal{D}_{\bm{\theta}_{3}}^{R}(\hat{\bm{I}}^{l-1}\oplus\bm{F}_{\text{novel}}),l\in\left\{1\cdots L\right\},$ (7)
where $\hat{\bm{I}}^{l}\in\mathbb{R}^{H\times W\times 3}$ and $\bm{F}_{\text{novel}}\in\mathbb{R}^{H\times W\times 32}$. Similar to the Gaussian attribute estimation, the loss function for supervising the final output is a combination of the MSE loss and the SSIM loss between the refined image $\hat{\bm{I}}^{L}$ and the ground truth $\bm{I}^{\text{gt}}$ as follows:
@@
Table 1: Comparison of computational and temporal costs for different attention mechanisms. CVA refers to the cross-view attention in MVSplat, EA denotes the epipolar attention in pixelSplat, and w16, w32, w64 represent the window sizes of 16, 32, and 64, respectively, in EVA. G.M. indicates the GPU memory usage.
Input Size 2$\times$64$\times$128$\times$128 2$\times$64$\times$256$\times$256 2$\times$32$\times$256$\times$256
Module Time G.M. Time G.M. Time G.M.
CVA 35.3ms 3.72GB 30.4ms 35.44GB 26.3ms 31.77GB
EA 58.3ms 15.19GB 19.3ms 59.14GB 16.9ms 58.01GB
EVA (w16) 7.23ms 0.92GB 1.77ms 2.14GB 1.43ms 1.37GB
EVA (w32) 6.53ms 0.92GB 1.49ms 2.14GB 1.16ms 1.37GB
EVA (w64) 6.31ms 0.92GB 1.39ms 2.14GB 1.06ms 1.37GB
@@ 3.5 Attribute Regularization
Existing feed-forward 3D Gaussian Splatting methods, such as GPS-Gaussian [36] and MVSplat [4], typically assume that the position of 3D Gaussians correspond directly to the estimated depth values in their respective views. However, this assumption does not always hold, as 3D Gaussians are not explicitly constrained to lie on the surface of objects. Moreover, the unprojection process in these methods is performed independently for each view using an orthogonal projection framework. This approach lacks explicit scene understanding and cross-view consistency, as it does not incorporate mechanisms for cross-view matching or global scene optimization during the standard training procedure.
To address these limitations and improve reconstruction quality, we introduce a regularization term to enhance the training process. Specifically, we propose an anchor loss to regularize the scales and opacities of Gaussians, ensuring consistency between the geometry of predicted depth maps and the 3D Gaussian positions, as proved in Appendix D. This loss also aligns the Gaussians from different views to force their locations to the same landmark. We leverage MediaPipe [16] to annotate human facial landmarks and compute the anchor loss to regularize the 3D landmark Gaussian scales, opacities, and positions as follows:
$\mathcal{L}_{\text{anchor}}=$ $\lambda_{\text{opacity}}\sum_{i=1}^{N}||\bm{O}_{i}\text{log}(\bm{O}_{i})||_{1}+\lambda_{\text{scale}}\sum_{i=1}^{N}||\bm{S}_{i}||_{2}+$
$\sum_{i,j\in\mathbb{V}}\sum_{m_{i}\in\mathbb{M}_{i},m_{j}\in\mathbb{M}_{j}}\max\left\{Dist,t\right\},$ (9)
where $Dist$ represents the distance of human facial landmarks in 3D space with
$Dist=||\Pi^{-1}(\bm{m}_{i},\bm{P}_{i}(\bm{m}_{i}))-\Pi^{-1}(\bm{m}_{j},\bm{P}_{j}(\bm{m}_{j}))||_{2},$
$\{\mathbb{M}_{i}\}_{i=1}^{n}$ denotes the collection of all landmarks on the 2D image plane, $\mathbb{V}$ denotes the collection of source views, and $\Pi^{-1}$ represents the process of reprojection from 2D image to 3D space. In addition, a tolerance factor $t$ is introduced to mitigate the errors in the estimated landmarks.
By integrating the loss functions from the three stages, i.e., $\mathcal{L}_{\text{depth}},\mathcal{L}_{\text{render}},\mathcal{L}_{\text{refine}}$, and the proposed regularization term $\mathcal{L}_{\text{anchor}}$, the overall training loss for EVA-Gaussian is:
$\mathcal{L}_{\text{EVA}}=\mathcal{L}_{\text{depth}}+\lambda_{1}\mathcal{L}_{\text{render}}+\lambda_{2}\mathcal{L}_{\text{refine}}+\lambda_{3}\mathcal{L}_{\text{anchor}},$ (10)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weighting factors.
Since the 3D Gaussian position and attribute estimation stages can be executed within tens of milliseconds, and the feature refinement stage is lightweight, taking less than ten milliseconds, EVA-Gaussian is capable of rapidly reconstructing 3D human subjects from a collection of RGB images and rendering novel views in a real-time manner.
@@ 4 Experiments
Table 2: Comparison with feed-forward 3D reconstruction methods at a resolution of 256$\times$256. Inference speeds are reported in the last column. Better results are marked in a deeper color.
$\Delta$=45° THuman2.0 THumansit Speed
@@
Table 3: Comparison of feed-forward methods under different camera angle settings at a resolution of $1024\times 1024$. Better results are marked in a deeper color. GPS-Gaussian fails to work effectively when $\Delta=90^{\circ}$, as it is unable to meet its rectification requirement.
@@
Figure 4: Qualitative comparison on THuman2.0 and THumansit. EVA-Gaussian achieves superior novel view rendering quality under diverse camera settings. Additional visualization results are provided in Appendix C.
4.1 Experiment Setup
Implementation details. Our EVA-Gaussian is trained on 1024$\times$1024 pixel images across multiple training views using a single NVIDIA A800 GPU for 100K iterations with the AdamW [15] optimizer, unless otherwise specified. For the 3D Gaussian position estimation stage, it is first pretrained under the supervision of ground truth depth maps. Baselines are trained using their publicly available code. More implementation details are provided in Appendix A.
Datasets. We conduct experiments on two open-source human body datasets: THuman2.0 [33] and THumanSit [34]. THuman2.0 contains 526 unique human models with their corresponding SMPL parameters, among which 100 individuals are randomly selected for our evaluation. The THumanSit dataset has a similar structure, containing 72 human models with around 60 poses for each, and we randomly choose 5 individuals with all poses for our evaluation.
Metrics. We report results on commonly used metrics: PSNR, SSIM [28], and LPIPS [35], computed over the entire image, as some methods may produce artifacts outside the human bounding box [11, 36]. We also include the inference time to demonstrate the real-time reconstruction capability of our method.
Table 4: Comparison with GPS-Gaussian under different camera numbers. Results in bold represent the best performance. EVA-Gaussian achieves SOTA performance across various metrics, primarily due to the multi-view consistency enabled by our proposed EVA module.
@@ 4.1 Experiment Setup
4.2 Stereo Reconstruction
Comparison with state-of-the-art feed-forward reconstruction methods. We first compare our approach against state-of-the-art (SOTA) feed-forward reconstruction methods, including ENeRF [11], pixelSplat [2], MVSplat [4], MVSGaussian [13], and GPS-Gaussian [36]. All experiments are conducted in a stereo-view setting, where the angle between the two camera views $\Delta=45^{\circ}$. The attention modules in the scene reconstruction methods [2, 13, 4] are inefficient in their utilization of GPU memory, limiting their ability to train effectively at a high resolution of 1024$\times$1024. Therefore, we also conduct a fair comparison of all methods at a resolution of 256$\times$256. The quantitative results presented in Table 2 demonstrate that EVA-Gaussian achieves the best novel view quality in terms of PSNR, SSIM, and LPIPS, while maintaining the second-fastest inference speed.
4.3 Multi-view Reconstruction
We conduct experiments under multi-view settings to evaluate EVA-Gaussian’s capability ti handle more than two input images. Table 4 presents the quantitative results, where our method demonstrates a performance gain with more than 1.5 dB improvement over the baseline. While the performance of GPS-Gaussian drops significantly due to the mismatch among multiple inferences, our method maintains high performance, thanks to the cross-view consistency ensured by our proposed EVA module.
4.4 Generalizability Validation
We validate the generalizability of EVA-Gaussian by taking out-of-domain images as input. It is observed that EVA-Gaussian is inherently superior in generalizing to out-of-distribution human identities and postures, primarily due to the strong data processing ability of its attention modules in EVA-Gaussian. This allows EVA-Gaussian to maintain consistent performance when provided with sufficient data. As shown in Table 5, given that THumansit contains significantly more human models (over 4,000) than THuman2.0 (526 models), EVA-Gaussian shows a greater performance improvement (+2.11 dB in PSNR) compared to GPS-Gaussian (+1.97 dB in PSNR). This conclusion is further supported by the evaluation on THumansit, where models trained on THuman2.0 experience a performance decline due to limited data availability. Despite this, EVA-Gaussian still outperforms GPS-Gaussian, achieving a performance gain of 0.41 dB in PSNR. In addition, Fig. 5 provides visual evidence of EVA-Gaussian’s robustness to out-of-domain data, since we have explicitly introduced inductive bias in the EVA module mitigates cross-domain performance degradation. Additional validation results on a real-world dataset are provided in Appendix B.
Table 5: Quantitative results of cross-domain validation. EVA-Gaussian consistently outperforms GPS-Gaussian.
Method THumansit $\rightarrow$ THuman2.0 THuman2.0 $\rightarrow$ THumansit
[…]
GPS-Gaussian 29.33 0.9733 0.0325 20.86 0.9243 0.0872
EVA-Gaussian 30.40 0.9751 0.0321 21.27 0.9275 0.0876
Figure 5: Visualization of cross-domain evaluation results for EVA-Gaussian. The left side displays the rendered results generated by EVA-Gaussian trained on the THuman2.0 dataset and evaluated on the THumansit dataset, while the right side shows the rendered results from EVA-Gaussian trained on the THumansit dataset and evaluated on the THuman2.0 dataset.
4.5 Ablation Study
We conduct a detailed ablation study on THuman2.0 in a stereo-view setting, where the angle between the two views $\Delta=45^{\circ}$, as shown in Table 6 and Fig. 6. We gradually incorporate the EVA module, feature refinement module, and anchor loss to evaluate their individual contributions. Removing the EVA module results in significant degradation across all metrics, as the network struggles to perform multi-view 3D Gaussian geometry prediction. When feature refinement is excluded, artifacts appear in critical areas, such as the hands and feet. Moreover, the lack of anchor loss leads to unreliable geometry predictions, particularly in the facial region, which in turn degrades the performance across all metrics, with a notable impact on LPIPS.
@@ 4.5 Ablation Study
w/o EVA w/o f.r. w/o anchor loss Full model
@@ 5 Conclusion
In this paper, we introduce EVA-Gaussian, a novel real-time 3D human reconstruction pipeline that employs multi-view attention-based 3D Gaussian position estimation and comprehensive feature refinement. To ensure robust performance, the method is trained using both photometric loss and anchor loss. Quantitative and qualitative evaluations on benchmark datasets demonstrate that EVA-Gaussian achieves state-of-the-art performance while maintaining a competitive inference speed, particularly under sparse camera settings.